# Two-phase flows through porous media described by a Cahn–Hilliard–Brinkman model with dynamic boundary conditions

### v1

@@ 4   Analysis of the Cahn–Hilliard–Brinkman system with singular potentials
We are now dealing with the proof of the existence of weak solutions for singular potentials. Our strategy is to approximate the convex parts of the singular potentials $F$ and $G$ satisfying (S1) and (S2) by means of a Yosida regularization. In this way, the approximate potentials are regular and exhibit quadratic growth and we can thus use Theorem 2.4 and Theorem 2.5 to obtain suitable approximate solutions. We then derive uniform estimates with respect to the approximation parameter, and eventually pass to the limit. In the forthcoming analysis, the splitting $F^{\prime}=\beta+\pi$ and $G^{\prime}={\beta_{\Gamma}}+\pi_{\Gamma}$ from (S1) will be adopted.
4.1   Yosida regularization
As mentioned, we rely on a Yosida regularization to smooth the singular parts of the potentials $F$ and $G$. For any $\varepsilon\in(0,1)$, we approximate the maximal monotone graphs $\beta$ and ${\beta_{\Gamma}}$ by
@@ 4.1   Yosida regularization
Then, the condition in (2.26) implies that
@@
(see, e.g., [15, Appendix]), where $\kappa_{1}$ and $\kappa_{2}$ are the constants introduced in (2.26). Next, we define $F_{\varepsilon}:=\widehat{\beta_{\varepsilon}}+\widehat{\pi},$ $G_{\varepsilon}:=\widehat{\beta}_{\Gamma,\varepsilon}+\widehat{\pi}_{\Gamma},$ where
$\widehat{\beta_{\varepsilon}}(r):=\int_{0}^{r}{\beta_{\varepsilon}}(s)\,{\rm ds},\quad\widehat{\beta}_{\Gamma,\varepsilon}(r):=\int_{0}^{r}\beta_{\Gamma,\varepsilon}(s)\,{\rm ds},\quad r\in{\mathbb{R}}.$
It is well-known that for every $r\in{\mathbb{R}}$,
@@
$&\hbox{For every $M>0$ there exist $C_{M}>0$ and $\varepsilon_{M}\in(0,1)$ such that}\\ &\widehat{\beta_{\varepsilon}}(r)\geq M\,r^{2}-C_{M}\quad\hbox{for every $r\in{\mathbb{R}}$ and every $\varepsilon\in(0,\varepsilon_{M})$}\,.\end{aligned}$ (4.3)
This property is checked in detail in the paper [20, beginning of Section 3]. Obviously, as a consequence, a similar condition holds for $\widehat{\beta}_{\Gamma,\varepsilon}$ since (4.1) entails that
@@
thanks to ${\beta_{\varepsilon}}(0)=\beta_{\Gamma,\varepsilon}(0)=0$ and since $\widehat{\beta_{\varepsilon}}$ and $\widehat{\beta}_{\Gamma,\varepsilon}$ have the same sign. Due to their construction by the Yosida approximation, $\beta_{\varepsilon}$ and $\beta_{\Gamma,\varepsilon}$ are Lipschitz continuous and have at most linear growth. Hence, $\widehat{\beta}_{\varepsilon}$ and $\widehat{\beta}_{\Gamma,\varepsilon}$ have at most quadratic growth. Moreover, (4.3) and (4.4) along with the (at most) quadratic growth of $\widehat{\pi}$ and ${\widehat{\pi}}_{\Gamma}$ (cf. (S1)), imply that both $F_{\varepsilon}$ and $G_{\varepsilon}$ are bounded from below by negative constants independent of $\varepsilon$. We can thus assume, without loss of generality, that $F_{\varepsilon}$ and $G_{\varepsilon}$ are nonnegative (otherwise, we add the modulus of their lower bounds, respectively). This entails that the approximate potentials $F_{\varepsilon}$ and $G_{\varepsilon}$ satisfy assumption (R1) with $p=q=2$.
@@ 4.2   Uniform estimates
for all $t\in[0,T]$. Now, observe that
@@
since $(\varphi_{0},\psi_{0})\in{\cal V}_{K}$ satisfies (2.8) and (4.2a) holds. Hence, in view of (A2) and (A3) and thanks to (4.5) and the Poincaré inequality, it is not difficult to infer that
@@
$\Big{|}\int_{0}^{T}\mathopen{\langle}\partial_{t}\varphi_{\varepsilon},\zeta\mathclose{\rangle}_{V}\Big{|}$ $\leq C\int_{0}^{T}(\left\|\varphi_{\varepsilon}\right\|_{L^{4}(\Omega)}\left\|{\boldsymbol{v}}_{\varepsilon}\right\|_{{\bf{L}}^{4}(\Omega)}+M_{2}\left\|\nabla\mu_{\varepsilon}\right\|_{{\bf H}})\left\|\nabla\zeta\right\|_{{\bf H}}$
@@
To deal with the terms on the left-hand side of (4.2), we recall that due to assumption (2.8), $m_{0}$ and $m_{\Gamma 0}$ lie in the interior of the domains $D(\beta)$ and $D(\beta_{\Gamma})$, respectively. We can thus exploit a useful property (see, e.g., [54, Appendix, Prop. A.1] and/or the detailed proof given in [36, p. 908]), namely there exist positive constants $c_{1},c_{2}$ and a nonnegative constant $c_{3}$ such that
@@
For the integrals in the second line of (4.2), we employ Poincaré’s inequality in the bulk and on the surface to obtain that
@@
[…]
@@
$\left\|F^{\prime}_{\varepsilon}(\varphi_{\varepsilon})\right\|_{L^{2}(0,T;L^{1}(\Omega))}+\left\|G^{\prime}_{\varepsilon}(\psi_{\varepsilon})\right\|_{L^{2}(0,T;L^{1}(\Gamma))}\leq C,$
since $\pi$ and $\pi_{\Gamma}$ are Lipschitz continuous. Consequently, by testing (2.13d) first by $(1,0)$ and then by $(0,1)$, one easily realizes that
@@
whence, using (4.2) and Poincaré’s inequality once again, we infer that
@@
Hence, with the help of Poincaré’s inequality, we arrive at
@@
Similarly, multiplying (4.16) by $-\beta_{\Gamma,\varepsilon}(\psi_{\varepsilon})$, it is straightforward to deduce that
@@
$-\Delta\varphi_{\varepsilon}$ $=\mu_{\varepsilon}-\beta_{\varepsilon}(\varphi_{\varepsilon})-\pi(\varphi_{\varepsilon})$ in $\quad\text{in $\Omega$},$
$-\Delta_{\Gamma}\psi_{\varepsilon}+\partial_{\bf n}\varphi_{\varepsilon}$ $=\theta_{\varepsilon}-\beta_{\Gamma,\varepsilon}(\psi_{\varepsilon})-\pi_{\Gamma}(\psi_{\varepsilon})$ on $\quad\text{on $\Gamma$},$
@@
a.e. in $(0,T)$. Due to (4.2), (4.28), (4.31) and the Lipschitz continuity of $\pi$ and $\pi_{\Gamma}$, it is clear that the right-hand sides in the above system belong to $L^{2}(\Omega)$ and $L^{2}(\Gamma)$, respectively. Hence, applying regularity theory for elliptic problems with bulk-surface coupling (see [48, Theorem 3.3]), we deduce that the estimate
$\left\|\varphi_{\varepsilon}\right\|_{H^{2}(\Omega)}+\left\|\psi_{\varepsilon}\right\|_{H^{2}(\Gamma)}$
$\quad{}\leq C\Bigl{(}\left\|\mu_{\varepsilon}-\beta_{\varepsilon}(\varphi_{\varepsilon})-\pi(\varphi_{\varepsilon})\right\|_{H}+\left\|\theta_{\varepsilon}-\beta_{\Gamma,\varepsilon}(\psi_{\varepsilon})+\psi_{\varepsilon}-\pi_{\gamma}(\psi_{\varepsilon})\right\|_{{H}_{\Gamma}}\Bigl{)}$
@@
for almost all $t\in(0,T)$. On this basis, at this point we can use the standard trace theorem for the normal derivative concluding that, for some fixed 3/$23/2<s<$ 2 there is a positive constant $C_{s}$ such that
@@ 4.3   Passage to the limit and conclusion of the proof
[…]
strongly in C0([0,T];Hs(Ω)), and a.e. in $\quad\text{strongly in $C^{0}([0,T];H^{s}(\Omega))$, and {a.e.\ in~{}$Q$}},$
[…]
strongly in C0([0,T];Hs(Γ)), and a.e. on $\quad\text{strongly in $C^{0}([0,T];H^{s}(\Gamma))$, and {a.e.\ on~{}$\Sigma$}},$
@@
$\varphi_{\varepsilon}$ $\to\varphi^{*}$ weakly in $\quad\text{weakly in $L^{2}(0,T;H^{2}(\Omega))$},$
$\psi_{\varepsilon}$ $\to\psi^{*}$ weakly in $\quad\text{weakly in $L^{2}(0,T;H^{2}(\Gamma))$}.$
@@
(see, e.g., [9, Prop. 1.1, p. 42]). Due to the aforementioned strong convergences of $\varphi_{\varepsilon}$ and $\psi_{\varepsilon}$, it is straightforward to check that condition (iii) of Definition 2.8 is fulfilled. Moreover, condition (iv) of Definition 2.8 can be established by proceeding analogously as in Subection 3.1.4.
@@
Proof of (a). Since ${\cal V}_{K}$ and ${\cal H}$ are separable Hilbert spaces with compact embedding ${\cal V}_{K}{\hookrightarrow}{\cal H}$ and continuous embedding ${\cal H}{\hookrightarrow}{\cal V}^{*}_{K}$, the assertion follows directly from the Lions–Magenes lemma (see, e.g., [60, Chapter III, Lemma 1.2]).
Proof of (b). We first fix $u$ and $v$ as arbitrary representatives of their respective equivalence class. Recall that due to (a), we have $u\in C^{0}([0,T];H)$ and $v\in C^{0}([0,T];{H}_{\Gamma})$. We can thus extend the functions $u$ and $v$ onto $[-T,0]$ by defining $u(t)$ and $v(t)$ by reflection for all $t<0$.
@@
$u_{k}$ $\to u$ strongly in $\quad\text{strongly in $L^{2}(0,T;H^{3}(\Omega))$},$ (A.3)
$v_{k}$ $\to v$ strongly in $\quad\text{strongly in $L^{2}(0,T;H^{3}(\Gamma))$},$ (A.4)
$(u_{k},v_{k})$ $\to(u,v)$ strongly in $\quad\text{strongly in $L^{2}(0,T;{\cal V}_{K})$},$ (A.5)
$(\partial_{t}u_{k},\partial_{t}v_{k})$ $\to(\partial_{t}u,\partial_{t}v)$ strongly in $\quad\text{strongly in $L^{2}(0,T;{\cal V}^{*})$}.$ (A.6)
@@
Since $(u_{k},v_{k})\to(u,v)$ strongly in $L^{2}\big{(}0,T;(H^{3}(\Omega)\times H^{3}(\Gamma))\big{)}$, we can fix $s\in[0,t]$ such that $(u_{k},v_{k})(s)\to(u,v)(s)$ strongly in $H^{3}(\Omega)\times H^{3}(\Gamma)$. Recalling the convergences (A.3)–(A.6), we thus infer that the right-hand side in (Proof.) tends to zero as $j,k\to\infty$. Consequently, $(\nabla u_{k})_{k\in{\mathbb{N}}}$ is a Cauchy sequence in $C^{0}([0,T];{\bf{H}})$ and $(\nabla v_{k})_{k\in{\mathbb{N}}}$ is a Cauchy sequence in $C^{0}([0,T];\mathbf{H}_{\Gamma})$. We thus conclude
$\nabla u_{k}$ $\to\nabla u$ strongly in $\quad\text{strongly in $C^{0}([0,T];{\bf{H}})$},$ (A.10)

### v2

@@ 4  Analysis of the Cahn–Hilliard–Brinkman system with singular potentials
We are now dealing with the proof of the existence of weak solutions for singular potentials. Our strategy is to approximate the convex parts of the singular potentials $F$ and $G$ satisfying (S1) and (S2) by means of a Moreau–Yosida regularization. In this way, the approximate potentials are regular and exhibit quadratic growth and we can thus use Theorem 2.6 and Theorem 2.7 to obtain suitable approximate solutions. We then derive uniform estimates with respect to the approximation parameter, and eventually pass to the limit. In the forthcoming analysis, the splitting $F^{\prime}=\beta+\pi$ and $G^{\prime}={\beta_{\Gamma}}+\pi_{\Gamma}$ from (S1) will be adopted.
4.1  Yosida regularizations
As mentioned, we rely on a Yosida regularization acting on the graphs $\beta$ and $\beta_{\Gamma}$}. For any $\varepsilon\in(0,1)$, we approximate the maximal monotone graphs $\beta$ and $\beta_{\Gamma}$} by
@@ 4.1  Yosida regularizations
It is well-known that $\beta_{\varepsilon}$ and $\beta_{\Gamma,\varepsilon}$ are single-valued and can be interpreted as monotone functions $\beta_{\varepsilon}:{\mathbb{R}}\to{\mathbb{R}}$ and $\beta_{\Gamma,\varepsilon}:{\mathbb{R}}\to{\mathbb{R}}$. Moreover, the condition in (2.30) implies that
@@
(see, e.g., [20, Appendix]), where $\kappa_{1}$ and $\kappa_{2}$ are the constants introduced in (2.30). Next, we define $F_{\varepsilon}:=\widehat{\beta_{\varepsilon}}+\widehat{\pi}$, $G_{\varepsilon}:=\widehat{\beta}_{\Gamma,\varepsilon}+\widehat{\pi}_{\Gamma},$ where
$\widehat{\beta_{\varepsilon}}(r):=\int_{0}^{r}{\beta_{\varepsilon}}(s)\,{\rm ds},\quad\widehat{\beta}_{\Gamma,\varepsilon}(r):=\int_{0}^{r}\beta_{\Gamma,\varepsilon}(s)\,{\rm ds},\quad r\in{\mathbb{R}}$
are actually the Moreau–Yosida regularizations of the singular parts $\widehat{\beta}$ and $\widehat{\beta}_{\Gamma}$ of the potentials $F$ and $G$. Now, it is well-known that for every $r\in{\mathbb{R}}$,
@@
[…]
This property is checked in detail in the paper [25, beginning of Section 3]. Obviously, as a consequence, a similar condition holds for $\widehat{\beta}_{\Gamma,\varepsilon}$ since (4.1) entails that
@@
thanks to ${\beta_{\varepsilon}}(0)=\beta_{\Gamma,\varepsilon}(0)=0$ and since $\widehat{\beta_{\varepsilon}}$ and $\widehat{\beta}_{\Gamma,\varepsilon}$ have the same sign. Due to their construction by the Yosida approximation, $\beta_{\varepsilon}$ and $\beta_{\Gamma,\varepsilon}$ are Lipschitz continuous and have at most linear growth. Hence, $\widehat{\beta}_{\varepsilon}$ and $\widehat{\beta}_{\Gamma,\varepsilon}$ have at most quadratic growth.
In the following, we assume that $\varepsilon\in(0,\varepsilon_{1})$, where $\varepsilon_{1}$ is given by (4.3) with $M=1$. Then, (4.3) with $M=1$ and (4.4) along with the (at most) quadratic growth of $\widehat{\pi}$ and ${\widehat{\pi}}_{\Gamma}$ (cf. (S1)), imply that both $F_{\varepsilon}$ and $G_{\varepsilon}$ are bounded from below by negative constants independent of $\varepsilon\in(0,\varepsilon_{1})$. We can thus assume, without loss of generality, that $F_{\varepsilon}$ and $G_{\varepsilon}$ are nonnegative (otherwise, we add the modulus of their lower bounds to $\widehat{\pi}$ or $\widehat{\pi}_{\Gamma}$, repsectively). In summary, this entails that the approximate potentials $F_{\varepsilon}$ and $G_{\varepsilon}$ satisfy assumption (R1) with $p=q=2$.
@@ 4.2  Uniform estimates
for all $t\in[0,T]$. Now, recalling that $F_{\varepsilon}$ and $G_{\varepsilon}$ satisfy assumption (R1) with $p=q=2$, we observe that
@@
since $(\varphi_{0},\psi_{0})\in{\cal V}_{K}$ satisfies (2.31) and (4.2a) holds. At this point, we recall that the potentials $F_{\varepsilon}$ and $G_{\varepsilon}$ were assumed (without loss of generality) to be nonnegative. Hence, in view of (A2) and (A3) and thanks to (4.5) and the Poincaré–Wirtinger inequality in $\Omega$ and Poincaré’s inequality on $\Gamma$ (see Lemma 2.2), it is not difficult to infer that
@@
$\Big{|}\int_{0}^{T}\mathopen{\langle}\partial_{t}\varphi_{\varepsilon},\zeta\mathclose{\rangle}_{V}\Big{|}$ $\leq C\int_{0}^{T}(\left\|\varphi_{\varepsilon}\right\|_{L^{6}(\Omega)}\left\|{\boldsymbol{v}}_{\varepsilon}\right\|_{{\bf{L}}^{3}(\Omega)}+M_{2}\left\|\nabla\mu_{\varepsilon}\right\|_{{\bf H}})\left\|\nabla\zeta\right\|_{{\bf H}}$
@@
To deal with the terms on the left-hand side of (4.2), we recall that due to assumption (2.31), $m_{0}$ and $m_{\Gamma 0}$ lie in the interior of the domains $D(\beta)$ and $D(\beta_{\Gamma})$, respectively. We can thus exploit a useful property (see, e.g., [68, Appendix, Prop. A.1] and/or the detailed proof given in [46, p. 908]), namely there exist positive constants $c_{1},c_{2}$ and a nonnegative constant $c_{3}$ such that
@@
For the integrals in the second line of (4.2), we employ Hölder’s inequality along with the Poincaré–Wirtinger inequality in $\Omega$ and Poincaré’s inequality on $\Gamma$ (see Lemma 2.2) to obtain that
@@
$\mu_{\varepsilon}$ $=-\Delta\varphi_{\varepsilon}+F^{\prime}_{\varepsilon}(\varphi_{\varepsilon})$ $\text{a.e. in $Q$},$ (4.15)
@@
$\left\|\pi(\varphi_{\varepsilon})\right\|_{L^{2}(0,T;L^{1}(\Omega))}+\left\|\pi_{\Gamma}(\psi_{\varepsilon})\right\|_{L^{2}(0,T;L^{1}(\Gamma))}\leq C$
since $\pi$ and $\pi_{\Gamma}$ are Lipschitz continuous. In combination with (4.20), this entails
$\left\|F^{\prime}_{\varepsilon}(\varphi_{\varepsilon})\right\|_{L^{2}(0,T;L^{1}(\Omega))}+\left\|G^{\prime}_{\varepsilon}(\psi_{\varepsilon})\right\|_{L^{2}(0,T;L^{1}(\Gamma))}\leq C.$
Consequently, by testing (2.16d) first by $(1,0)$ and then by $(0,1)$, one easily realizes that
@@
whence, using (4.2), the Poincaré–Wirtinger inequality in $\Omega$ and Poincaré’s inequality on $\Gamma$ (see Lemma 2.2), we infer that
@@
Hence, with the help of the Poincaré–Wirtinger inequality in $\Omega$ and Poincaré’s inequality on $\Gamma$ (see Lemma 2.2), we arrive at
@@
Similarly, multiplying (4.16) by $\beta_{\Gamma,\varepsilon}(\psi_{\varepsilon})$, it is straightforward to deduce that
@@
$-\Delta\varphi_{\varepsilon}$ $=\mu_{\varepsilon}-\beta_{\varepsilon}(\varphi_{\varepsilon})-\pi(\varphi_{\varepsilon})$ $\quad\text{in $\Omega$},$
$-\Delta_{\Gamma}\psi_{\varepsilon}+\partial_{\bf n}\varphi_{\varepsilon}$ $=\theta_{\varepsilon}-\beta_{\Gamma,\varepsilon}(\psi_{\varepsilon})-\pi_{\Gamma}(\psi_{\varepsilon})$ $\quad\text{on $\Gamma$},$
@@
a.e. in $(0,T)$. Due to (4.2), (4.28), (4.31) and the Lipschitz continuity of $\pi$ and $\pi_{\Gamma}$, it is clear that the right-hand sides in the above system belong to $L^{2}(\Omega)$ and $L^{2}(\Gamma)$, respectively. Hence, applying regularity theory for elliptic problems with bulk-surface coupling (see [61, Theorem 3.3]), we deduce that the estimate
$\begin{aligned} &\left\|\varphi_{\varepsilon}\right\|_{H^{2}(\Omega)}+\left\|\psi_{\varepsilon}\right\|_{H^{2}(\Gamma)}\\ &\quad{}\leq C\Bigl{(}\left\|\mu_{\varepsilon}-\beta_{\varepsilon}(\varphi_{\varepsilon})-\pi(\varphi_{\varepsilon})\right\|_{H}+\left\|\theta_{\varepsilon}-\beta_{\Gamma,\varepsilon}(\psi_{\varepsilon})+\psi_{\varepsilon}-\pi_{\gamma}(\psi_{\varepsilon})\right\|_{{H}_{\Gamma}}\Bigl{)}\end{aligned}$
@@
for almost all $t\in(0,T)$. On this basis, at this point we can use the standard trace theorem for the normal derivative concluding that, for some fixed $3/2<s<2$ there is a positive constant $C_{s}$ such that
@@ 4.3  Passage to the limit and conclusion of the proof
$\varphi_{\varepsilon}$ $\to\varphi^{*}$ $\text{weakly-${}^{*}$ in $L^{\infty}(0,T;V)$, weakly in $H^{1}(0,T;V^{*})$},$
$\quad\text{strongly in $C^{0}([0,T];H^{s}(\Omega))$, and {a.e.\ in~$Q$}},$
$\psi_{\varepsilon}$ $\to\psi^{*}$ $\text{weakly-${}^{*}$ in $L^{\infty}(0,T;{V}_{\Gamma})$, weakly in $H^{1}(0,T;{V}_{\Gamma}^{*})$},$
$\quad\text{strongly in $C^{0}([0,T];H^{s}(\Gamma))$, and {a.e.\ on~$\Sigma$}},$
@@
$\varphi_{\varepsilon}$ $\to\varphi^{*}$ $\quad\text{weakly in $L^{2}(0,T;H^{2}(\Omega))$},$
$\psi_{\varepsilon}$ $\to\psi^{*}$ $\quad\text{weakly in $L^{2}(0,T;H^{2}(\Gamma))$}.$
@@
(see, e.g., [12, Prop. 1.1, p. 42]). Due to the aforementioned strong convergences of $\varphi_{\varepsilon}$ and $\psi_{\varepsilon}$, it is straightforward to check that condition (iii) of Definition 2.9 is fulfilled. Moreover, condition (iv) of Definition 2.9 can be established by proceeding analogously as in Subection 3.1.4.
@@
[…]
Proof of (b). We first fix $u$ and $v$ as arbitrary representatives of their respective equivalence class. Recall that due to (a), we have $u\in C^{0}([0,T];H)$ and $v\in C^{0}([0,T];{H}_{\Gamma})$. We can thus extend the functions $u$ and $v$ onto [T,0][-T,0] by defining $u(t)$ and $v(t)$ by reflection for all $t<0$.
@@
$u_{k}$ $\to u$ $\quad\text{strongly in $L^{2}(0,T;H^{3}(\Omega))$},$ (A.3)
$v_{k}$ $\to v$ $\quad\text{strongly in $L^{2}(0,T;H^{3}(\Gamma))$},$ (A.4)
$(u_{k},v_{k})$ $\to(u,v)$ $\quad\text{strongly in $L^{2}(0,T;{\cal V}_{K})$},$ (A.5)
$(\partial_{t}u_{k},\partial_{t}v_{k})$ $\to(\partial_{t}u,\partial_{t}v)$ $\quad\text{strongly in $L^{2}(0,T;{\cal V}^{*})$}.$ (A.6)
@@
Since $(u_{k},v_{k})\to(u,v)$ strongly in $L^{2}\big{(}0,T;(H^{3}(\Omega)\times H^{3}(\Gamma))\big{)}$, we can fix $s\in[0,t]$ such that $(u_{k},v_{k})(s)\to(u,v)(s)$ strongly in $H^{3}(\Omega)\times H^{3}(\Gamma)$. Recalling the convergences (A.3)–(A.6), we thus infer that the right-hand side in (Appendix: Some calculus for bulk-surface function spaces) tends to zero as $j,k\to\infty$. Consequently, $(\nabla u_{k})_{k\in{\mathbb{N}}}$ is a Cauchy sequence in $C^{0}([0,T];{\bf{H}})$ and $(\nabla v_{k})_{k\in{\mathbb{N}}}$ is a Cauchy sequence in $C^{0}([0,T];\mathbf{H}_{\Gamma})$. We thus conclude
$\nabla u_{k}$ $\to\nabla u$ $\quad\text{strongly in $C^{0}([0,T];{\bf{H}})$},$ (A.10)